OpenStudy (anonymous):

solve this radical equation square root1-3x=-7 I'm trying to graduate guys and girls and i have just a short amount of time to finish up my work so please help me.

OpenStudy (anonymous):

if your factoring?

OpenStudy (anonymous):

or solve for x

OpenStudy (anonymous):

yeah i suck at factoring

OpenStudy (anonymous):

and i am not good with solving for x

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

thanks sidneysider

OpenStudy (anonymous):

(x+1)^2=X^2+3 X^2+2x+1=x^2+3 X^2+2x+1-x^2-3=0 2x-2=0 x=1